OpenStudy (anonymous):

Find the center and radius of the circle x^2-6x+y^2=7

OpenStudy (anonymous):

(x-3)^2 - 9 + y^2 = 7 (x - 3)^2 + y^2 = 16 center is (3,0) and radius = 4

OpenStudy (anonymous):

Where did u get 0?

OpenStudy (anonymous):

y^2 = (y-0)^2

OpenStudy (anonymous):

i compared the equation to standard result (x-a)^2 + (y-b)^2 = r^2

OpenStudy (anonymous):

ohhh Ok thank YOU! life saver..can you take a look at my other unanswered question?

OpenStudy (anonymous):

ok